Calculate the pressure of hydrogen gas produced if 50 grams of Aluminum




reacts with 250 grams of Hydrochloric acid. The Hydrogen gas is collected




at 2.75 L vessel at 80.6 ℉.

1
Expert's answer
2021-11-24T02:07:02-0500

Al +6HCl "\to" 2AlCl3 + 3H2

50g 250g

P= nRT / V


Moles of All


=m / RFM


= 50/27

= 1.8519moles


Moles of HCl= m/ RFM


= 250 / 36.5


= 6.8493 moles


Moles of Al = 6.8493 / 6


= 1.1416moles


Moles of H2 =


( 3 × 6.8493 / 6 ) moles


n = 3.42465 moles

V= 2.75l

T=86.6 °F

If 32°F= 0 °C

"\therefore" 80.6 ° F


80.6 / 32 × 1


= 2.5188 ° C

T in "\kappa"

2.51888 + 273.15


=275.6688k

r = nRT / V

(3.42465 × 8.3142 × 275.

6688k ) / 2.75


= 2854.2472 atm